OpenStudy (anonymous):

Solve the system of linear equations below. y = -2x + 19 y = x + 7

OpenStudy (anonymous):

Substitute x+7 from second line into 1st line: x+7 = -2x + 19 : Simplify x+2x+7=19 3x+7=19 3x=12 x=4 (answer)